find the equation of the circle which passes through the origin and cuts the chords of length 4 and 6 on the +ve side of x-axix and y-axis

Grade:8

1 Answers

md salik
11 years ago

general eqn of circle is x2 + y2 +2gx + 2fy + c =0

since circle passes through origin

for (0,0)

c=0

since circle cuts chord of length 4 unit

for (4,0)

16 + 8g =0

g=-2

for 6 unit on y axis

for (0,6)

36 + 12f =0

f = -3

therefore eqn of circle is x2 + y2 -4x - 6y =0

is the required equation